James is 12 Years older than his brother Tim. Three years ago, James was three times as old as Tim. Find their presents ages

Answer by jorel1380(3719)

Let Tim be t. Then James is t+12. So:
t+12-3=3(t-3)
t+9=3t-9
2t=18
t=9
Tim is 9; James is 21. ☺☺☺☺